OpenStudy (anonymous):

There are two boxes. You have 10 black and 10 white chess pieces. How do you arrange the pieces in the two boxes, so that the probability to pick a white piece is the largest?

OpenStudy (anonymous):

Help please :(

OpenStudy (anonymous):

does it matter?

OpenStudy (anonymous):

Not sure, heh, that's what I am trying to find out

OpenStudy (anonymous):

are you picking a box and random, and then picking a piece at random from a box? question is unclear as written

OpenStudy (anonymous):

Yes, you pick a box at random and then a piece at random from the box.

OpenStudy (anonymous):

I agree that the question was a bit unclear

OpenStudy (anonymous):

if you put all white in one box, probability is \(\frac{1}{2}\) see what happens if you change it, i am willing to bet it makes no difference

OpenStudy (anonymous):

That was my thought too, but it seemed a little too easy

OpenStudy (anonymous):

for example 5 and 5 probability is \(\frac{1}{2}\times \frac{1}{2}+\frac{1}{2}\times \frac{1}{2}=\frac{1}{2}\)

OpenStudy (anonymous):

try it with 3 and 7 i gotta run

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{7}{10}+\frac{1}{2}\times \frac{3}{10}=\frac{1}{2}(\frac{3+7}{10})\] the same

OpenStudy (anonymous):

should give an easy proof

OpenStudy (anonymous):

Thank you :)

OpenStudy (zarkon):

try putting one white piece in a box and all the others in the second box

OpenStudy (anonymous):

can you explain the reasoning?

OpenStudy (anonymous):

would be very helpful

OpenStudy (anonymous):

I got the same answer from another person, but I don't understand it, it seems counterlogical to me

OpenStudy (zarkon):

why...the fists box there is a 100% chance of getting a white and the second there is a 9/19 change so prob of white is \[\frac{1}{2}\times 1+\frac{1}{2}\frac{9}{19}=\frac{14}{19}\]

OpenStudy (zarkon):

which is a lot bigger than 1/2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

But why is there a 9/19 chance in the second? Sorry, I suck at this

OpenStudy (zarkon):

there are only 19 pieces in the second box (one in the first) of those 9 are white

OpenStudy (anonymous):

Ohhhhhhh, just realized

OpenStudy (anonymous):

I've gone around thinking thinking there had to be 10 in each

OpenStudy (anonymous):

Hah, thanks alot :)

OpenStudy (zarkon):

would be a boring problem then ;)

OpenStudy (anonymous):

Indeed :)

OpenStudy (anonymous):

My intuition messed things up for me